OpenStudy (anonymous):

I would like some help doing these two problems. If you could, please explain the problems, that would be even better. I need these simplified, but just don't see how it can be done. 1/(1+x/y) x+1/x^2+2x-8 - x/4x-8

OpenStudy (anonymous):

Calmat?

OpenStudy (anonymous):

To simplify the first one, think of it as two fractions, one on top of the other. The bottom term is 1+x/y. If we get the common denominator as y, \[1+\frac{ x }{ y }=\frac{ y+x }{ y }\] Now we have to divide 1 by this denominator. To do so just requires us to multiply 1 by the reciprocal of the fraction. So, \[1\div \frac{ y+x }{ y }=1\times \frac{ y }{ y+x }=\frac{ y }{ y+x }\]. The second one requires a little more work. \[\frac{ x+1 }{ x ^{2}+2x-8 }-\frac{ x }{ 4x-8 }=\frac{ x+1 }{ (x+4)(x-2) }-\frac{ x }{ 4(x-2)}\] Now we have to find a common denominator. Both denominators have a common factor of (x-2), so that is their greatest common factor. Multiply it by the factors which they don't have in common. This leads us to an LCD of : 4(x-2)(x+4). So, we must convert each denominator to this LCD. This leads us to the following: \[\frac{ 4(x+1) }{4(x+4)(x-2) }-\frac{ x(x+4) }{4(x-2)(x+4) }\] Now, multiply each numerator out. This leads us to: \[\frac{ 4x+4-x ^{2}-4x }{4(x+4)(x-2)}\] Combine like terms in the numerator. \[\frac{ -x ^{2} +4}{ 4(x+4)(x-2) }\] Finally, factor out a -1 from the numerator and factor the result as the difference of squares. \[\frac{ -1(x-2)(x+2) }{ 4(x+4)(x-2) }\] From here, the common factor x-2 may be reduced. Leaving your final answer as: \[\frac{ -(x+2) }{ 4(x+4) }\]

OpenStudy (anonymous):

Sorry, took me a while to type that second one out.

OpenStudy (anonymous):

No problem, it is perfectly all right. Your explanations are superb! Thank you so very much for taking the time to help mt out!

OpenStudy (anonymous):

no problem. You are most welcome.

OpenStudy (anonymous):

These are spelled out so easily that even I can understand them, ha. The only class I ever really have an issue with is Algebra 2. This helps me out more than you know!

OpenStudy (anonymous):

Good, I am glad it helps you.

OpenStudy (anonymous):

Thanks again!